# Vikhr: Constructing a State-of-the-art Bilingual Open-Source Instruction-Following Large Language Model for Russian

### v5

@@ 4.3 Quality of Generated Text
Previous results demonstrate the performance of our model in NLU tasks. However, the performance in NLU does not always reflect the performance in text generation, as the former is based on classification and ranking capabilities of models.
@@ 4.4 Ablation Studies
We conducted several ablation studies to demonstrate effects of various features of our model translation pipeline that allowed Vikhr to achieve high performance.
@@
is illustrated in Table 10. The results show that the KL regularization introduced in our pipeline slightly reduces perplexity and substantially increases the model performance in the ruXNLI task. This shows that continued pre-training on its own might deteriorate LLM reasoning capabilities, and proper regularization helps preventing the catastrophic forgetting.
@@ 5 Conclusion
We have presented Vikhr – a new state-of-the-art open-source instruction-following bilingual LLM oriented on the Russian language. To create Vikhr, we developed a comprehensive pipeline for adapting English-oriented LLMs to other languages. The pipeline includes the adaptation of the tokenizer vocabulary, continued pre-training of the entire model, and instruction tuning. We have also constructed a new dataset for instruction tuning by expanding the Saiga dataset with automatically translated and cleaned English instruction datasets. Our extensive work enabled Vikhr to outperform the known baselines while maintaining computational efficiency.
We hope that the developed cross-lingual adaption pipeline and the published models will foster the research on LLMs and enhance the diversity of languages incorporated into research agendas.
@@
We do not introduce additional restrictions to the usage of our models. However, the users must comply with the license of the base model and instruction datasets.
We do not implement RLHF / DPO fine-tuning of Vikhr due to the lack of the resources for human annotation. We expect further performance improvements from these techniques.

### v6

@@ 4.3 Ablation Studies
We conducted several ablation studies to demonstrate the effects of various features of our model translation pipeline.
@@
is illustrated in Table 9. The results show that the KL regularization introduced in our pipeline slightly reduces perplexity and substantially increases the model performance in the ruXNLI task. This shows that continued pre-training on its own might deteriorate LLM reasoning capabilities, and proper regularization helps prevent the catastrophic forgetting.
@@ 5 Conclusion
We have presented Vikhr – a new open-source instruction-following bilingual LLM oriented on the Russian language. To create Vikhr, we developed a comprehensive pipeline for adapting English-oriented LLMs to other languages. The pipeline includes the adaptation of the tokenizer vocabulary, continued pre-training of the entire model, and instruction tuning. We have also constructed a new dataset for instruction tuning by expanding the Saiga dataset with automatically translated and cleaned English instruction datasets. Our extensive work enabled Vikhr to achieve high quality of generation while maintaining computational efficiency.
We hope that the developed cross-lingual adaptation pipeline and the published models will foster the research on LLMs and enhance the diversity of languages incorporated into research agendas.
@@
We do not introduce additional restrictions on the usage of our models. However, the users must comply with the license of the base model and instruction datasets.
We do not implement RLHF / DPO fine-tuning of Vikhr due to the lack of resources for human annotation. We expect further performance improvements from these techniques.